200,000 barrels per day (BPD) fractionation tower costs $1,200,000 when the appropriate cost index was 360.0. Estimate the cost of a 300,000 BPD fractionation tower when the index is 600.0. The exponent for the cost-capacity equation is 0.67
Answer #1

Cost of 200000 Capacity at index of 600 = 1200000 * 600 / 360 = 2000000

Cost of 300000 Capacity at index of 600 = 2000000 * (300000 / 200000)^0.67

= 2624285.85